# Video-Language Critic: Transferable Reward Functions for Language-Conditioned Robotics

### v2

@@ Abstract
Natural language is often the easiest and most convenient modality for humans to specify tasks for robots. However, learning to ground language to behavior typically requires impractical amounts of diverse, language-annotated demonstrations collected on each target robot. In this work, we aim to separate the problem of what to accomplish from how to accomplish it, as the former can benefit from substantial amounts of external observation-only data, and only the latter depends on a specific robot embodiment. To this end, we propose Video-Language Critic, a reward model that can be trained on readily available cross-embodiment data using contrastive learning and a temporal ranking objective, and use it to score behavior traces from a separate actor. When trained on Open X-Embodiment data, our reward model enables 2x more sample-efficient policy training on Meta-World tasks than a sparse reward only, despite a significant domain gap. Using in-domain data but in a challenging task generalization setting on Meta-World, we further demonstrate more sample-efficient training than is possible with prior language-conditioned reward models that are either trained with binary classification, use static images, or do not leverage the temporal information present in video data.
Keywords: Vision-Language Models, Reinforcement Learning, Reward Learning
@@ 1 Introduction
Advances in natural language processing and vision-language representations have enabled a significant increase in the scalability and generalization abilities of learned control policies for robotics. Methods involving large architectures, such as Transformers [1], and internet-scale pretraining have transferred well to both high-level [2, 3] and low-level [4, 5, 6] robotic control. Natural language has many desirable features as a modality for specifying tasks. Unlike structured, hand-designed task sets, natural language is unrestricted and open-domain. Moreover, prompts can be specified as precisely or vaguely as appropriate. While goal images, demonstration videos, or goal states more broadly, have been considered as an alternative open-domain task definition modality [7, 8, 9], they typically have to specify irrelevant environment details, such as the background. Furthermore, language readily supports task definitions with novel combinations of actions, objects and their attributes, as well as subtask sequencing, in a way that facilitates the policy’s understanding of unseen tasks.
@@
Most prior work has proposed to learn language-conditioned policies end-to-end, i.e., directly predicting an action in the robot’s action space given the current state and task description. However, this has several downsides: first, fitting large models on the full problem requires a significant amount of high-quality demonstration data from the target robot domain. Second, the resulting policy depends entirely on the specific robot instance, observation space, and controller type (e.g., joint or task space control) and does not easily transfer to other settings. Moreover, much of the prior work addresses vision-language grounding in robotics purely with imitation learning [4, 5, 6, 10], without attempting to discriminate between low-quality and expert demonstrations. As a result, the resulting policies are inherently limited by the skills of the demonstrators, and no novel solutions can be discovered through planning or interaction. This line of work overlooks performance gains that could be obtained by converting the language prompts to a scalar reward function. Manually defining a well-specified dense reward function to communicate task success is typically laborious and error-prone, and must be repeated for each task. To make progress towards a general-purpose robotic system that can learn human-level skills both in terms of quality (dexterity, robustness) and variety of skills, we argue these systems will need to be able to critique their own behavior, by learning reward functions at scale.
@@
We address this problem by learning a foundation video-language-conditioned reward model, i.e., a critic that evaluates the progress (in the form of a video) of a task, given as a human-language instruction, and assigns a reward based on how close the robot is to completing the task. By leveraging large cross-task pretraining data, which may come from a variety of robots, our Video-Language Critic (VLC), can learn to score the alignment between a textual description and task execution regardless of the specific robot embodiment. Our experimental evaluation on Meta-World [11] manipulation tasks shows that VLC can learn useful general-purpose reward functions not only from in-domain, but also out-of-domain data (Open X-Embodiment [10]) collected from different robot embodiments. In Section 4, we show that VLC 1) accelerates the training of a wide range of manipulation tasks and 2) enables zero-shot learning on unseen tasks, when combined with a sparse task completion signal.
@@
Figure 1: Overview: Our similarity function, $S_{\theta}$, is trained using video-caption pairs $(v_{1:T}^{i},c^{i})$. The visual encoder (ViT, [12]) is applied separately to each video frame $v_{t}$ to produce a sequence of image features, which are appended to the caption embedding produced by the text encoder. The temporal aggregator then predicts a similarity score for each time step $t$ of the video. We use cross-entropy and sequence ranking objectives to encourage the predicted scores to be high for matching video-caption pairs and low for mismatching pairs, and to monotonously increase over a successful execution.
@@
Recent work in language-conditioned rewards for robotics has used either binary classification [13, 14, 15], contrastive vision-language alignment [16, 17, 18] or reconstruction [19] objectives. However, they have not fully leveraged temporal ordering of frames to encourage increasing scores over successful episodes. In contrast, VLC provides a dense reward evaluating in-episode progress, learned through contrastive ranking. The advantages and contributions of our approach are as follows:
@@
  • Through maximizing the learned reward, our policies can improve over sub-optimal demonstrations, by executing the task faster or by finding better solutions.
@@ 2 Related Work
Many prior works have aimed to connect language instructions and vision-based observations in robotics [4, 5, 6, 20, 21] and in video games [22], mostly through large-scale demonstrations [4, 5, 20, 22] or pretraining [6, 21]. However, the majority of approaches have considered imitation-based objectives only, without ranking existing trajectories or attempting to outperform prior data. We instead propose to learn a state-value function from cross-domain offline behavior, which can be optimized using either online, offline or model-based policy training.
@@
Pretrained vision-language representations [23] have been adapted to a wide range of downstream tasks [6, 21]. Shridhar et al. [6] propose to augment pretrained CLIP [23] with Transporter nets [24] to handle fine-grained spatial awareness required for precise manipulation. Xiao et al. [25] train a CLIP-like contrastive embedding space from crowd-sourced language annotations for trajectories from the robot. We draw inspiration from these works, but instead define an embodiment-agnostic, language-conditioned reward function, which supports improvement over demonstration data.
@@
Our work is related to video retrieval as we seek to move beyond image-language correspondence and match task descriptions with history-aware state representations. As learning representations across time is computationally expensive, many prior works have proposed to start from pretrained image-language representations and aggregate them over time, while fine-tuning the aggregation function’s weights on video retrieval [26, 27, 28]. Unlike in video retrieval, we aim to not only assign high alignment scores to full videos, but provide smoothly increasing reward over the whole video to indicate task progress.
@@
Several works have proposed to infer the reward function of a task using examples of expert behavior, and to train an RL policy to optimize this reward [29]. Most relevantly to our setting, a line of prior inverse RL methods considers the case where the observed behavior is not annotated with actions and may come from different action and observation spaces altogether, typically a human demonstrator [7, 8, 9, 13, 14, 15, 30, 31, 32]. Many of these works use either a goal image [8, 9, 32] or a demonstration video [7, 31] rather than language conditioning, and some are only applicable for data from a single task at a time [30, 32]. Moreover, handling multi-task reward learning with an additional task identifier state variable, as done by Chen et al. [7], requires a predefined grouping into a discrete set of tasks. By contrast, our use of language to define tasks enables a more subtle and composable task space.
@@
Although a few prior works have used unrestricted natural language to define rewards for robotic manipulation tasks using either binary classification [13, 14, 15] or contrastive vision-language alignment [22, 18, 16, 17], these methods have not fully leveraged the temporal ordering of frames in their objective to encourage increasing scores over a successful episode. We instead propose to explicitly learn increasing rewards for partial trajectories making progress towards solving the task. Moreover, most prior methods use only a single image or a pair of images, whereas we consider the full episode to better represent partially observable tasks. Moreover, many prior works only considered data from the actor’s own observation space [14, 15, 22], whereas we propose to learn from cross-embodiment data allowing zero-shot transfer to a robot with different morphology, kinematics, and visual appearance.
Specifically, a few prior works [22, 33, 34] have explored CLIP-based models for learning vision-language reward functions. However, these works have been limited to image features of the current time step [33, 34] or a snapshot of the most recent history [22]. We instead propose to score behavior at the time series level by comparing the task description and a full video trajectory. Despite its name, RoboCLIP [16] (proposed to solve a similar problem to our work) is not based on the CLIP architecture or the pretrained representations unlike our method, but the S3D architecture [35], and mainly shares the InfoNCE loss with CLIP.
@@ 3 Video-Language Critic
We propose to learn language-conditioned robotic manipulation by first training an embodiment-agnostic reward function on video-caption pairs, and then using the learned reward model to guide the autonomous training of a robot-specific policy. To serve as a useful reward signal for downstream policy learning, the learned function should accurately represent the intended task, while providing enough signal to the agent to enable efficient learning [36, 37, 38, 39]. It should exhibit at least two key properties: accuracy and temporal smoothness. Making progress in the specified task should be rewarded with positive feedback with as little delay as possible, i.e., the function should smoothly increase over a successful execution. In fact, the problem of optimal reward shaping is equivalent to learning the value function for the optimal policy [39], suggesting that an optimal densely shaped reward should monotonously increase over a successful demonstration (assuming the reward we ultimately wish to maximize corresponds to sparse goal reaching). Moreover, the end-of-episode scores for successful trajectories should exceed those of incomplete or failed executions: classification accuracy between successes and failures should be high. With these desiderata, we formulate Video-Language Critic, a language-conditioned reward model trained with cross-entropy and sequential ranking objectives to encourage progressively increasing scores over a successful video’s duration.
@@
Our approach is motivated by the success of contrastive image-language pretraining and the wide applicability of pretrained CLIP [23] encoders as foundation models. The problem of comparing observed behavior to a desired task description is analogous to the setting of CLIP; however, we extend the contrastive learning approach to scoring videos. Compared to a single image, using sequences of frames sampled across the full trajectory increases the generality of our reward function, and could allow it to handle non-Markovian (i.e., history-dependent) tasks. Such tasks might involve partial observability, repetitive or circular movements, or be described relative to an earlier state; even simple object displacement tasks may fall in this category.
@@
We define video and text encoder networks similar to CLIP4Clip used for video retrieval [26], the task of finding videos within an existing dataset that most closely match a given textual query. The general architecture is shown in Fig. 1. First, each video frame is processed with an image encoder network while the video caption is processed with a text encoder, both initialized with CLIP in order to benefit from its large-scale vision-language pretraining. Luo et al. [26] tested different aggregation strategies for reasoning over the resulting sequence of image features. In video retrieval, averaging image features over time was found to be sufficient, and no performance benefit could be obtained with an attention-based aggregation. While video retrieval shares similarities with our setting, task progress evaluation requires a much more nuanced understanding of temporal dynamics: for one, reversing the video should typically result in a very different reward value.
To support this, it is necessary for the temporal aggregation function to process the input frames as an ordered sequence, which in the case of Transformer aggregators is achieved using position embeddings. Furthermore, we find that embedding both modalities independently of each other and comparing with a cosine similarity, as done by CLIP as well as CLIP4Clip’s sequence Transformer aggregation, causes the resulting video representation to lose too much of the information relevant to task completion. Thus, instead of using cosine similarity, we train a single temporal aggregation Transformer to directly output a similarity score based on the concatenated textual and image features. Each encoder’s weights are fine-tuned and the aggregation function is trained from scratch for the video-text matching task. We refer to the full architecture, consisting of both the encoder networks and the aggregation Transformer, as a similarity function $S_{\theta}$, parameterized by $\theta$.
@@
As training signal, we wish to leverage weak supervision from video-level captions without known spatial or temporal extent. We use a contrastive objective function to encourage each caption to better match its corresponding video than other videos, and vice versa. We therefore train the similarity prediction network $S_{\theta}$ with symmetric cross-entropy as done by Radford et al. [23], i.e., with the mean of text-to-video and video-to-text cross-entropy terms, for video-caption pairs $(v^{i},c^{i}),i=1..N$:
@@
Video inputs also contain implicit information about the relative ranking of states, which is not leveraged in prior reward learning approaches [15, 22, 16, 18, 19, 17]. We propose to learn from this temporal signal by extending the cross-entropy objective with a sequential ranking term. Each subsequent state in a successful trajectory should, in general, have higher value for completing the task than its predecessors, which the reward function should reflect. Our total loss then becomes:

### v3

@@ Abstract
Natural language is often the easiest and most convenient modality for humans to specify tasks for robots. However, learning to ground language to behavior typically requires impractical amounts of diverse, language-annotated demonstrations collected on each target robot. In this work, we aim to separate the problem of what to accomplish from how to accomplish it, as the former can benefit from substantial amounts of external observation-only data, and only the latter depends on a specific robot embodiment. To this end, we propose Video-Language Critic, a reward model that can be trained on readily available cross-embodiment data using contrastive learning and a temporal ranking objective, and use it to score behavior traces from a separate actor. When trained on Open X-Embodiment data, our reward model enables 2x more sample-efficient policy training on Meta-World tasks than a sparse reward only, despite a significant domain gap. Using in-domain data but in a challenging task generalization setting on Meta-World, we further demonstrate more sample-efficient training than is possible with prior language-conditioned reward models that are either trained with binary classification, use static images, or do not leverage the temporal information present in video data. [^1] Source code and supplementary videos are available on the project website: https://sites.google.com/view/video-language-critic.
@@ 1 Introduction
Advances in natural language processing and vision-language representations have enabled a significant increase in the scalability and generalization abilities of learned control policies for robotics. Methods involving large architectures, such as Transformers (Vaswani et al., 2017), and internet-scale pretraining have transferred well to both high-level (Liang et al., 2022; Vemprala et al., 2023) and low-level (Brohan et al., 2022; Lynch et al., 2022; Shridhar et al., 2022) robotic control. Natural language has many desirable features as a modality for specifying tasks. Unlike structured, hand-designed task sets, natural language is unrestricted and open-domain. Moreover, prompts can be specified as precisely or vaguely as appropriate. While goal images, demonstration videos, or goal states more broadly, have been considered as an alternative open-domain task definition modality (Chen et al., 2021; Alakuijala et al., 2023; Ma et al., 2023b), they typically have to specify irrelevant environment details, such as the background. Furthermore, language readily supports task definitions with novel combinations of actions, objects and their attributes, as well as subtask sequencing, in a way that facilitates the policy’s understanding of unseen tasks.
@@
Most prior work has proposed to learn language-conditioned policies end-to-end, i.e., directly predicting an action in the robot’s action space given the current state and task description. However, this has several downsides: first, fitting large models on the full problem requires a significant amount of high-quality demonstration data from the target robot domain. Second, the resulting policy depends entirely on the specific robot instance, observation space, and controller type (e.g., joint or task space control) and does not easily transfer to other settings. Moreover, much of the prior work addresses vision-language grounding in robotics purely with imitation learning (Brohan et al., 2022; Lynch et al., 2022; Shridhar et al., 2022; Open X-Embodiment Collaboration et al., 2023), without attempting to discriminate between low-quality and expert demonstrations. As a result, the resulting policies are inherently limited by the skills of the demonstrators, and no novel solutions can be discovered through planning or interaction. This line of work overlooks performance gains that could be obtained by converting the language prompts to a scalar reward function. Manually defining a well-specified dense reward function to communicate task success is typically laborious and error-prone, and must be repeated for each task. To make progress towards a general-purpose robotic system that can learn human-level skills both in terms of quality (dexterity, robustness) and variety of skills, we argue these systems will need to be able to critique their own behavior, by learning reward functions at scale.
@@
We address this problem by learning a foundation video-language-conditioned reward model, i.e., a critic that evaluates the progress (in the form of a video) of a task, given as a human-language instruction, and assigns a reward based on how close the robot is to completing the task. By leveraging large cross-task pretraining data, which may come from a variety of robots, our Video-Language Critic (VLC) can learn to score the alignment between a textual description and task execution regardless of the specific robot embodiment. Our experimental evaluation on Meta-World (Yu et al., 2019) manipulation tasks shows that VLC can learn useful general-purpose reward functions not only from in-domain, but also out-of-domain data (from Open X-Embodiment Collaboration et al. (2023)) collected from different robot embodiments. In Section 4, we show that VLC 1) accelerates the training of a wide range of manipulation tasks and 2) enables zero-shot learning on unseen tasks, when combined with a sparse task completion signal.
@@
Figure 1: Overview: Our similarity function, $S_{\theta}$, is trained using video-caption pairs $(v_{1:T}^{i},c^{i})$. The visual encoder (ViT; Dosovitskiy et al. (2021)) is applied separately to each video frame $v_{t}$ to produce a sequence of image features, which are appended to the caption embedding produced by the text encoder. The temporal aggregator then predicts a similarity score for each time step $t$ of the video. We use cross-entropy and sequence ranking objectives to encourage the predicted scores to be high for matching video-caption pairs and low for mismatching pairs, and to monotonically increase over a successful execution.
@@
Recent work in language-conditioned rewards for robotics has used either binary classification (Shao et al., 2021; Silva et al., 2021; Nair et al., 2022a), contrastive vision-language alignment (Nair et al., 2022b; Ma et al., 2023a; Sontakke et al., 2023) or reconstruction (Karamcheti et al., 2023) objectives. However, they have not fully leveraged temporal ordering of frames to encourage increasing scores over successful episodes. In contrast, VLC provides a dense reward evaluating in-episode progress, learned through contrastive ranking. The advantages and contributions of our approach are as follows:
@@
  • Through maximizing the learned reward, our policies can improve over suboptimal demonstrations, by executing the task faster or by finding better solutions.
@@ 2 Related Work
Many prior works have aimed to connect language instructions and vision-based observations in robotics (Lynch & Sermanet, 2020; Brohan et al., 2022; Lynch et al., 2022; Shridhar et al., 2022; Guhur et al., 2023) and in video games (Fan et al., 2022), mostly through large-scale demonstrations (Lynch & Sermanet, 2020; Brohan et al., 2022; Fan et al., 2022; Lynch et al., 2022) or pretraining (Shridhar et al., 2022; Guhur et al., 2023). However, the majority of approaches have considered imitation-based objectives only, without scoring the quality of existing trajectories or attempting to outperform prior data. We instead propose to learn a state-value function from cross-domain offline behavior, which can be optimized using either online, offline or model-based policy training.
@@
Pretrained vision-language representations (Radford et al., 2021) have been adapted to a wide range of downstream tasks (Shridhar et al., 2022; Guhur et al., 2023). Shridhar et al. (2022) propose to augment pretrained CLIP (Radford et al., 2021) with Transporter nets (Zeng et al., 2021) to handle fine-grained spatial awareness required for precise manipulation. Xiao et al. (2022) train a CLIP-like contrastive embedding space from crowd-sourced language annotations for trajectories from the robot. We draw inspiration from these works, but instead define an embodiment-agnostic, language-conditioned reward function, which supports improvement over demonstration data.
@@
Our work is related to video retrieval as we seek to move beyond image-language correspondence and match task descriptions with history-aware state representations. As learning representations across time is computationally expensive, many prior works have proposed to start from pretrained image-language representations and aggregate them over time, while fine-tuning the aggregation function’s weights on video retrieval (Bain et al., 2022; Luo et al., 2022; Lu et al., 2023). Unlike in video retrieval, we aim to not only assign high alignment scores to full videos, but provide smoothly increasing reward over the whole video to indicate task progress.
@@
Several works have proposed to infer the reward function of a task using examples of expert behavior, and to train an RL policy to optimize this reward (Russell, 1998). Most relevantly to our setting, a line of prior inverse RL methods considers the case where the observed behavior is not annotated with actions and may come from different action and observation spaces altogether, typically a human demonstrator (Sermanet et al., 2018; Schmeckpeper et al., 2020; Chen et al., 2021; Shao et al., 2021; Silva et al., 2021; Nair et al., 2022a; Zakka et al., 2022; Alakuijala et al., 2023; Ma et al., 2023b). Many of these works use either a goal image (Zakka et al., 2022; Alakuijala et al., 2023; Ma et al., 2023b) or a demonstration video (Sermanet et al., 2018; Chen et al., 2021) rather than language conditioning, and some are only applicable for data from a single task at a time (Schmeckpeper et al., 2020; Zakka et al., 2022). Moreover, handling multi-task reward learning with an additional task identifier state variable, as done by Chen et al. (2021), requires a predefined grouping into a discrete set of tasks. By contrast, our use of language to define tasks enables a more subtle and composable task space.
@@
Although a few prior works have used unrestricted natural language to define rewards for robotic manipulation tasks using either binary classification (Shao et al., 2021; Silva et al., 2021; Nair et al., 2022a) or contrastive vision-language alignment (Fan et al., 2022; Nair et al., 2022b; Ma et al., 2023a; Sontakke et al., 2023), these methods have not fully leveraged the temporal ordering of frames in their objective to encourage increasing scores over a successful episode. We instead propose to explicitly learn increasing rewards for partial trajectories making progress towards solving the task. Moreover, most prior methods use only a single image or a pair of images, whereas we consider the full episode to better represent partially observable tasks. Moreover, many prior works only considered data from the actor’s own observation space (Silva et al., 2021; Fan et al., 2022; Nair et al., 2022a), whereas we propose to learn from cross-embodiment data allowing zero-shot transfer to a robot with different morphology, kinematics, and visual appearance.
Specifically, a few prior works (Fan et al., 2022; Baumli et al., 2023; Rocamonde et al., 2023) have explored CLIP-based models for learning vision-language reward functions. However, these works have been limited to image features of the current time step (Baumli et al., 2023; Rocamonde et al., 2023) or a snapshot of the most recent history (Fan et al., 2022). We instead propose to score behavior at the time series level by comparing the task description and a full video trajectory. Despite its name, RoboCLIP (Sontakke et al., 2023) (proposed to solve a similar problem to our work) is not based on the CLIP architecture or the pretrained representations unlike our method, but the S3D architecture (Xie et al., 2018), and mainly shares the InfoNCE loss with CLIP.
@@ 3 Video-Language Critic
We propose to learn language-conditioned robotic manipulation by first training an embodiment-agnostic reward function on video-caption pairs, and then using the learned reward model to guide the autonomous training of a robot-specific policy. To serve as a useful reward signal for downstream policy learning, the learned function should accurately represent the intended task, while providing enough signal to the agent to enable efficient learning (Ackley & Littman, 1992; Singh et al., 2009; 2010; Sorg, 2011). It should exhibit at least two key properties: accuracy and temporal smoothness. Making progress in the specified task should be rewarded with positive feedback with as little delay as possible, i.e., the function should smoothly increase over a successful execution. In fact, the problem of optimal reward shaping is equivalent to learning the value function for the optimal policy (Sorg, 2011), suggesting that an optimal densely shaped reward should monotonically increase over a successful demonstration (assuming the reward we ultimately wish to maximize corresponds to sparse goal reaching). Moreover, the end-of-episode scores for successful trajectories should exceed those of incomplete or failed executions: classification accuracy between successes and failures should be high. With these desiderata, we formulate Video-Language Critic, a language-conditioned reward model trained with cross-entropy and sequential ranking objectives to encourage progressively increasing scores over a successful video’s duration.
@@
Our approach is motivated by the success of contrastive image-language pretraining and the wide applicability of pretrained CLIP (Radford et al., 2021) encoders as foundation models. The problem of comparing observed behavior to a desired task description is analogous to the setting of CLIP; however, we extend the contrastive learning approach to scoring videos. Compared to a single image, using sequences of frames sampled across the full trajectory increases the generality of our reward function, and could allow it to handle non-Markovian (i.e., history-dependent) tasks. Such tasks might involve partial observability, repetitive or circular movements, or be described relative to an earlier state; even simple object displacement tasks may fall in this category.
@@
We define video and text encoder networks similar to CLIP4Clip used for video retrieval (Luo et al., 2022), the task of finding videos within an existing dataset that most closely match a given textual query. The general architecture is shown in Fig. 1. First, each video frame is processed with an image encoder network while the video caption is processed with a text encoder, both initialized with CLIP in order to benefit from its large-scale vision-language pretraining. Luo et al. (2022) tested different aggregation strategies for reasoning over the resulting sequence of image features. In video retrieval, averaging image features over time was found to be sufficient, and no performance benefit could be obtained with an attention-based aggregation. While video retrieval shares similarities with our setting, task progress evaluation requires a much more nuanced understanding of temporal dynamics: for one, reversing the video should typically result in a very different reward value.
To support this, it is necessary for the temporal aggregation function to process the input frames as an ordered sequence, which in the case of Transformer aggregators is achieved using position embeddings. Furthermore, we find that embedding both modalities independently of each other and comparing with a cosine similarity, as done by CLIP as well as CLIP4Clip’s sequence Transformer aggregation, causes the resulting video representation to lose too much of the information relevant to task completion. Thus, instead of using cosine similarity, we train a single temporal aggregation Transformer to directly output a similarity score based on the concatenated textual and image features. Our architecture is compared with the mean pooling and sequence Transformer variants in Figure 5. The text and the visual encoders’ weights are fine-tuned and the aggregation function is trained from scratch for the video-text matching task. We refer to the full architecture, consisting of both the encoder networks and the aggregation Transformer, as a similarity function $S_{\theta}$, parameterized by $\theta$.
@@
As training signal, we wish to leverage weak supervision from video-level captions without known spatial or temporal extent. We use a contrastive objective function to encourage each caption to better match its corresponding video than other videos, and vice versa. We therefore train the similarity prediction network $S_{\theta}$ with symmetric cross-entropy as done by Radford et al. (2021), i.e., with the mean of text-to-video and video-to-text cross-entropy terms, for video-caption pairs $(v^{i},c^{i}),i=1..N$:
@@
Video inputs also contain implicit information about the relative ranking of states, which is not leveraged in prior reward learning approaches (Fan et al., 2022; Nair et al., 2022a; b; Karamcheti et al., 2023; Ma et al., 2023a; Sontakke et al., 2023). [^2] While Ma et al. (2023a) use order information, they do not use relative ranking: consecutive frames are used bidirectionally as negatives in their contrastive objective, with their representations pushed apart from each other. This encourages the frames’ representations to vary gradually over the video time span but does not leverage any prior on the later state having a higher value. We propose to learn from this temporal signal by extending the cross-entropy objective with a sequential ranking term. Each subsequent state in a successful trajectory should, in general, have higher value for completing the task than its predecessors, which the reward function should reflect. Our total loss then becomes: